# Superuniversal Statistics with Topological Origins for non-Hermitian Scattering Singularities

### v1

@@ Abstract
Vortex singularities and their speckle patterns in random superpositions of waves are an inevitable consequence of destructive interference and are consequently generic and ubiquitous. Singularities are topologically stable, meaning they persist under small perturbations and can only be removed via pairwise annihilation. They have found application in optical, acoustic, elastic, and fluid wave fields, among others. We generalize the concept of singularity speckle patterns to arbitrary two dimensional parameter spaces and any complex scalar function that describes wave phenomena involving complicated scattering. In scattering systems specifically, we are often concerned with singularities associated with complex zeros of the scattering matrix $S$. Some examples are Coherent Perfect Absorption (CPA), Reflectionless Scattering Modes (RSMs), Transmissionless Scattering Modes (TSMs), and $S$-matrix Exceptional Points (EPs). Experimentally, we find that all scattering singularities share a universal statistical property: the tail of the probability distribution function of a quantity that diverges only at a singularity has the form of a $-3$ power law. The heavy tail of the distribution provides an estimate for the likelihood of finding a given singularity in a generic scattering system. We use these universal statistical results to determine that homogeneous system loss is the most important parameter determining singularity density in a given parameter space. Finally, we discuss events where distinct singularities coincide in parameter space, which result in higher order singularities that have new applications. These higher order singularities are not topologically protected, and we do not find universal statistical properties for them. We support our empirical results from microwave experiments with Random Matrix Theory simulations and conclude that all results presented are found in all generic scattering systems.
@@ I Introduction
Topological defects in superpositions of complex plane waves where the amplitude vanishes were first seriously studied by Nye and Berry in 1974 [1]. They showed these phase dislocations, later called vortices or singularities, are due to interference effects, not dispersion. Vortices are present in any complex scalar field, and come with integer winding numbers which, with parametric evolution of the system, allows for pair-wise creation and annihilation of vortices with opposite winding numbers [2]. Due to the ubiquity of singularities, they appear and have been studied in both classical and quantum systems [3], from speckle patterns in disordered optics [4, 5] to nodes of wavefunctions [6, 7], superconductors [8, 9] and quantum weak measurements [10].
@@
In a smooth complex scalar field $\mathcal{S}$ which occupies a $d$-parameter space, there are an infinite number of $d-1$ dimensional structures that are topologically stable, meaning their existence is robust to small perturbations of the field. An example of such a structure is the contour of $\text{Re}[\mathcal{S}]=0$ ($\text{Im}[\mathcal{S}]=0$), although any scalar value that is within the bounds of $\text{Re}[\mathcal{S}]$ ($\text{Im}[\mathcal{S}]$) will have the same stability. The intersection of two distinct $d-1$ dimensional structures of the same field results in topologically stable $d-2$ dimensional structures. These structures are the vortex singularities that have drawn so much interest. The simplest example is $\mathcal{S}=0+i0$ as the intersection of $\text{Re}[\mathcal{S}]=0$ and $\text{Im}[\mathcal{S}]=0$, but the intersection of $\text{Re}[\mathcal{S}]=a$ and $\text{Im}[\mathcal{S}]=b$ is the complex zero of the field $\mathcal{S}-(a+ib)$ [11]. In this paper, we will use the term “fundamental singularity” to refer to a vortex of a complex scalar field, whatever that field may be.
@@
However, while two distinct $d-1$ structures of a single field can intersect to create a $d-2$ structure, the intersection of two $d-2$ structures of a single field results in vortex-antivortex annihilation, so there are no $d-3$ or beyond dimensional structures. This is equivalent to stating that a complex scalar field has two degrees of freedom, $d-1$ structures are constraints on one of those degrees of freedom while $d-2$ structures are constraints on both. It is impossible to constrain more degrees of freedom than $\mathcal{S}$ has. Hence without loss of generality, we can use the language of the $d=2$ case for simplicity, that being topologically protected points ($d-2$) and curves ($d-1$) [12]. In the ensuing paragraphs, any mention of points and curves is implicitly referring to “points in two dimensional parameter space” and “curves in two dimensional parameter space”.
@@
A network of such points in two dimensions result in what are known as “speckle patterns” [13, 14, 15]. We reiterate the pervasiveness of singularities by referencing the use and study of speckle patterns for metrology [16, 17, 18], velocimetry [19, 20], non-destructive testing [21, 22], imaging techniques such as Digital Image Correlation [23, 24] and Speckle Contrast Imaging [25, 26], and more. There is also a rich literature about the statistical properties of speckle patterns [27, 28, 29, 5]. In this paper, the concept of speckle patterns is generalized to any complex scalar field in arbitrary parameter space. Singularities as curves in three dimensional parameter space can be seen in Fig. 2 of Ref. [15], and Appendix B.
In this paper we specifically consider the singularities of the scattering $S$-matrix of a non-Hermitian scattering cavity that is open to the outside world through $M$ asymptotic scattering channels. The scattering matrix relates ingoing $\psi_{in}$ and outgoing $\psi_{out}$ monochromatic waves by |$\ket{\psi_{out}}=S\ket{\psi_{in}}$, and can be related to the effective Hamiltonian of the closed cavity through the Heidelberg approach [30, 31, 32, 33, 34, 35]. We focus on the $S$-matrix because it is experimentally accessible in many scattering platforms (including acoustics, particle physics, quantum transport, non-destructive testing, electromagnetic waves, etc.) and contains a depth of information about the closed system and its scattering processes. Related to the $S$-matrix are the impedance $Z$, admittance $Y$, and Wigner reaction $K$ matrices [36] that have similar properties and show the same singularity statistics.
@@
One method through which information about a system can be extracted from the $S$-matrix is analysis of complex time delays. Complex time delays are energy (or equivalently frequency) derivatives of functions of the $S$-matrix (such as $\text{det}S$ or $S_{xx}$) and can be directly related to time and center-frequency shifts of time domain pulses propagating through the system [37, 38]. Time delays have previously been used to determine the locations of poles and zeros of $S$-matrix in the complex frequency plane [32, 39, 40, 41, 42, 43] which are the related to the eigenvalues of the system’s effective Hamiltonian [44]. Another established application of time delay is identifying optimal wave configurations with high sensitivity, for use in both cavity and wavefront shaping [45, 46, 47, 48].
@@
Recently it was found that the probability density functions (PDF) of every complex time delay quantity [40, 43, 49, 50, 51] in non-Hermitian systems has the same universal $-3$ power law tail [52, 53]. Since complex time delays diverge at singularities of their associated scattering parameter [48, 54], the tail of the PDFs give a measure of the abundance of these phenomena. In this paper we will go further to show that the PDF of any quantity that diverges solely at a scattering singularity has the same $-3$ power law tail independent of the function, singularity, or system parameters.
@@ I.1 Fundamental Singularities
One singularity that has drawn a lot of interest is known as anti-lasing or alternatively as Coherent Perfect Absorption (CPA). This is a phenomena characterized by all the energy injected into a system being completely absorbed, with no reflection or transmission [55, 39, 56, 57, 58, 48, 59, 60]. CPA has been found in optics [61, 62, 63], acoustics [64, 65, 66], heat transfer [67], and quantum single photon systems [68, 69, 70]. In microwave scattering, the signature of CPA is an eigenvalue of the $S$-matrix being zero, so $\text{det}S$ is the complex scalar function who’s zero is the enabling condition for CPA. Alongside enabling CPA, $\text{det}S=0+i0$ also turns a two channel system into a tunable robust splitter [71].
@@
Figure 1: Speckle pattern of scattering singularities in $d=2$ parameter space. The experimental S-matrix data comes from a tetrahedral microwave graph. (a-b) Magnitude of $\text{det}S$, red lines mark locations of $\text{Re}[\text{det}S]=0$ while black lines mark locations of $\text{Im}[\text{det}S]=0$. White symbols highlight intersections which correspond to $\text{det}S=0+i0$ which are CPAs. (c-d) Phase of $\text{det}S$, black circles show locations of phase windings which align exactly with the white symbols in (a-b). Arrows on the circles in (d) show the two nearby CPAs have opposite winding number.
Another singularity of interest is an Exceptional Point degeneracy of order $P$ (EP-P) where $P\geq 2$ eigenvalues and eigenvectors of a non-Hermitian operator become degenerate [72, 73, 74, 75, 76, 77]. Similar to CPAs, EPs have been explored in many fields such as optics [78, 79, 80, 81], microwave cavities [82, 83, 84], exciton-polaritons [85], acoustics [86], waveguides [87], electrical circuits [88, 89], etc. EP-2s are topologically protected as they are the zeros of the complex scalar field $\lambda_{j}-\lambda_{k}$ where $\lambda_{k}$ is the $k$th eigenvalue of an operator $\mathcal{H_{S}}$. Exceptional Points of order higher than $2$ are not topologically protected, as will be discussed later.
In this paper we specifically consider EPs of the $S$-matrix, which are generally unrelated to the Hamiltonian EPs, except for specially contrived cases such as in Ref. [90]. The EPs of the $S$-matrix are the same as those for the impedance $Z$, admittance $Y$, and Wigner reaction $K$ matrices [91]. For a unitary scattering matrix, the eigenvectors are always orthogonal, so there are no Exceptional Points, only Diabolic Points (DP). A subunitary scattering matrix can have orthogonal eigenvectors, but only as a special case, which makes the eigenvector orthogonality condition also of interest. We consider the two-eigenvector degeneracy and orthogonality both with one quantity, the eigenvector coalescence measure $|C_{jk}|$ where:
@@
is the normalized inner-product of two right eigenvectors of the scattering matrix such that $S\ket{R_{j}}=\lambda_{j}\ket{R_{j}}$[91]. The quantity $|C_{jk}|$ is bounded between 00 (|$\ket{R_{j}}$ and |$\ket{R_{k}}$ are orthogonal) and $1$ (|$\ket{R_{j}}$ and |$\ket{R_{k}}$ are degenerate, namely an EP-2). We find that for a system with Lorentz reciprocity, two-eigenvector orthogonality has the dimension of curves (similar to the red and black lines in Fig. 1), but if reciprocity does not hold then the orthogonality takes the dimensionality of points [91]. Going forward, only the coalescence of two eigenvectors will be considered in this paper, and the $jk$ notation will be suppressed. The eigenvector coalescence is related to the Petermann factor $K=\frac{1}{1-|C|^{2}}$ [92, 93, 94].
Other simple singularities include matrix-element zeros of the $S$-matrix. The complex zeros of $S_{xx}$ or $S_{xy}$ are known as Reflectionless Scattering Modes (RSMs) and Transmissionless Scattering Modes (TSMs) [95, 96, 97, 60, 98, 37, 54, 99, 100, 101]. In a system with broken Lorentz reciprocity, it is possible to have $S_{xy}=0$ or $S_{yx}=0$ independently of each other, and when only one equals zero there is Uni-Directional Transmission. Further we can define $\delta R_{xy}:=S_{xx}-S_{yy}$ and $\delta T_{xy}:=S_{xy}-S_{yx}$ as the reflection and transmission differences between channels $x$ and $y$. If $\delta R_{xy}=0+i0$, then the two channels have symmetric reflections, which is an interesting singularity when it occurs in complex scattering systems without any geometric symmetries. Similarly if $\delta T_{xy}=0+i0$, then transmission between the two channels is reciprocal and this singularity is present in systems with broken Lorentz reciprocity.
As stated before, the divergence of complex time delays is associated with scattering singularities: Wigner-Smith $\tau_{WS}=\frac{-i}{M}\frac{\partial}{\partial E}\text{log}[\text{det}S]$ with CPA, Reflection $\tau_{xx}=-i\frac{\partial}{\partial E}\text{log}[S_{xx}]$ with RSM, Transmission $\tau_{xy}=-i\frac{\partial}{\partial E}\text{log}[S_{xy}]$ with TSM, etc. Since complex time delay itself can also take the value $00+i$ 0, there are additional scattering singularities that can be identified with complex time delay. To see why the zero of complex time delay is interesting, we turn to recent work on interpreting complex time delay in the frequency domain as the time and center frequency shift of a pulse in the time domain [37, 38]. This implies that the zero of complex time delay corresponds to a minimal distortion point of the system, at which signals are only attenuated (or amplified), but not otherwise distorted after propagation.
There are a myriad of other scattering singularities that we could discuss, and a more expansive listing is provided in Appendix F.
@@ I.2 Higher Order Singularities
All the singularities given so far are topologically stable, being the zero of a single complex scalar field $\mathcal{S}$. In experimental systems, we can use embedded perturbers to create collisions of fundamental singularities of different fields ($\mathcal{S}$ and $\mathcal{S}^{\prime}$) at real frequencies in a given parameter space. These collisions are not topologically protected and a small fluctuation in the system can move the singularities apart. Despite their instability, such “higher order” singularities can have novel properties and further applications, and so should be of great interest.
One example of a higher order singularity is the $S$-matrix CPA-EP demonstrated in Ref. [91] to turn a complex, irregular tetrahedral graph into a robust 50:50 In-Phase/Quadrature microwave splitter. CPA-EPs have also been found in acoustic systems [90], and in the spectrum of the Hamiltonian [102, 103]. There are also higher order $S$-matrix EP-Ps which are intersections of exactly P11-1- 1 unique EP-2s. Because the eigenvalue detuning around an EP has a $\frac{1}{P}$ power law, it shows extreme sensitivity to small perturbations, leading to proposals of EP sensors [79, 80, 81, 104]. RSMs and TSMs can also be made degenerate, for example if both $S_{xx}$ and $S_{yy}$ equal zero simultaneously then the system has an RSM-2, sometimes called a RSM or RL EP [105] because the eigenvectors of a reflectionless operator become degenerate. RSM-2s have been used to create broadband perfect transmission [106]. Reciprocal systems have fundamental TSM-2s and can only have TSM-Ps of even order P.
@@
Again, there are many ways to combine fundamental singularities into higher order singularities, but none of these combinations are topologically stable and the more fundamental singularities included, the less probable such events become.

### v3

@@ Abstract
Vortex singularities in speckle patterns formed from random superpositions of waves are an inevitable consequence of destructive interference and are consequently generic and ubiquitous. Singularities are topologically stable, meaning they persist under small perturbations and can only be removed via pairwise annihilation. They have applications including sensing, imaging and energy transfer in multiple fields such as optics, acoustics, and elastic or fluid waves. We generalize the concept of singularity speckle patterns to arbitrary two dimensional parameter spaces and any complex scalar function that describes wave phenomena involving complicated scattering. In scattering systems specifically, we are often concerned with singularities associated with complex zeros of various functions of the scattering matrix $S$. Some examples are Coherent Perfect Absorption (CPA), Reflectionless Scattering Modes (RSMs), Transmissionless Scattering Modes (TSMs), and $S$-matrix Exceptional Points (EPs). Experimentally, we find that all scattering singularities share a universal statistical property: any quantity that diverges as a simple pole at a singularity, such as $\text{det}S$ at CPA, has a probability distribution function with a 3-3 power law tail. The heavy tail of the distribution provides an estimate for the likelihood of finding a given singularity in a generic scattering system. We use these universal statistical results to determine that homogeneous system loss is the most important parameter determining singularity density in a given parameter space of an absorptive scattering system. Finally, we discuss events where distinct singularities coincide in parameter space, which result in higher order singularities that have new applications beyond the capabilities of isolated singularities. These higher order singularities are not topologically protected, and we do not find universal statistical properties for them. We support our empirical results from microwave experiments with Random Matrix Theory simulations and conclude that the statistical results presented hold for all generic non-Hermitian scattering systems in which singularities can occur.
@@ I Introduction
Topological defects in superpositions of complex plane waves where the amplitude vanishes were first formally introduced by Nye and Berry in 1974 [Nye1974]. They showed these phase dislocations, later called vortices or singularities, are due to interference effects, not dispersion. Vortices are present in any complex scalar field, and come with integer winding numbers which, with parametric evolution of the system, allows for pair-wise creation and annihilation of vortices with opposite winding numbers [Neu1990]. Due to the ubiquity of singularities, they appear and have been studied in both classical and quantum systems [KT73, Bliokh2017], from speckle patterns in disordered optics [Freund1993, Zhang2007, Cao2022] to nodes of wavefunctions [Hirschfelder1974, Heckenberg1992], superconductors [Abrikosov57, Beasley79, Song2009, wang2025] and quantum weak measurements [Solli2004].
@@
In a smooth complex scalar field $\mathcal{S}$ which occupies a $d$-parameter space, there are an infinite number of d1d-1 dimensional structures that are topologically stable, meaning their existence is robust to small perturbations of the field. An example of such a structure is the contour of $\text{Re}[\mathcal{S}]=0$ ($\text{Im}[\mathcal{S}]=0$), although any scalar value that is within the bounds of $\text{Re}[\mathcal{S}]$ ($\text{Im}[\mathcal{S}]$) will have the same stability. The intersection of two distinct d1d-1 dimensional structures of the same field results in topologically stable d2d-2 dimensional structures. These structures are the vortex singularities that have drawn so much interest. The simplest example is $\mathcal{S}=0+i0$ as the intersection of $\text{Re}[\mathcal{S}]=0$ and $\text{Im}[\mathcal{S}]=0$, but the intersection of $\text{Re}[\mathcal{S}]=a$ and $\text{Im}[\mathcal{S}]=b$ is the complex zero of the field $\mathcal{S}-(a+ib)$ [Hajnal1987, Berry2000, DeAngelis2016]. In this paper, we will use the term “fundamental singularity” to refer to a vortex of a complex scalar field, whatever that field may be.
@@
However, while two distinct d1d-1 structures of a single field can intersect to create a d2d-2 structure, the intersection of two d2d-2 structures of a single field results in vortex-antivortex annihilation, so there are no d3d-3 or beyond dimensional structures. This is equivalent to stating that a complex scalar field has two degrees of freedom, d1d-1 structures are constraints on one of those degrees of freedom (such as constraining only $\text{Re}[\mathcal{S}]$), while d2d-2 structures are constraints on both (i.e. both $\text{Re}[\mathcal{S}]$ and $\text{Im}[\mathcal{S}]$). It is impossible to constrain more degrees of freedom than $\mathcal{S}$ has. Hence without loss of generality, we can use the language of the $d=2$ case for simplicity, that being topologically protected points (d2d-2) and curves (d1d-1) [Dennis2025]. In the rest of this paper, any mention of points and curves is implicitly referring to “points in two dimensional parameter space” and “curves in two dimensional parameter space”.
@@
A network of such points in two dimensions result in what are known as “speckle patterns” [Goodman1975, Goodman1976, Zhang2010]. We reiterate the pervasiveness of singularities by referencing the use and study of speckle patterns for metrology [Fujii1976, Stetson1987, Wang2006], velocimetry [Barker1977, Andres1999], non-destructive testing [YANG1995, BRUNO2000], imaging techniques such as Digital Image Correlation [Dong2017, Crammond2013] and Speckle Contrast Imaging [Briers2013, Heeman2019], and more. There is also a rich literature about the statistical properties of speckle patterns [Dainty1970, Wagner1983, Okamoto1995, Zhang2007]. In this paper, the concept of speckle patterns is generalized to any complex scalar field in arbitrary parameter space. Singularities as curves in three dimensional parameter space can be seen in Fig. 2 of Ref. [Zhang2010], and Appendix C.
[…]
@@
One method through which information about a system can be extracted from the $S$-matrix is analysis of complex time delays. Complex time delays are energy (or equivalently frequency) derivatives of functions of the $S$-matrix (such as $\text{det}S$ or $S_{xx}$) and can be directly related to time and center-frequency shifts of time domain pulses propagating through the system [Asano2016, giovannelli2025]. Time delays have previously been used to determine the locations of poles and zeros of $S$-matrix in the complex frequency plane [Fyodorov1997bTRI, Fyodorov2017, Fyodorov2019, Osman2020, Lei2021WS, Lei2022] which are the related to the eigenvalues of the system’s effective Hamiltonian [Osman2020, Lei2021WS, Lei2022, Ma2023]. Another established application of time delay is identifying optimal wave configurations with high sensitivity, for use in both cavity- and wavefront-shaping [Rotter2011, Gerardin2016, Horodynski2020, Hougne2021].
@@
Recently it was found that the probability density functions (PDF) of every complex time delay quantity [Fyodorov2019, Lei2022, Patel2023, Mao2023, Saalman2024] in non-Hermitian systems has the same universal 3-3 power law tail [Lei2021Stats, shaibe2025]. Since complex time delays diverge at singularities of their associated scattering parameter [Hougne2021, Huang2022], the tail of the PDFs give a measure of the abundance of these phenomena. In this paper we will go further to show that the PDF of any quantity that has a simple pole divergence solely at a scattering singularity has the same 3-3 power law tail independent of the function, singularity, or system parameters.
@@ I.1 Fundamental Singularities
One singularity that has drawn a lot of interest is known as anti-lasing or alternatively as Coherent Perfect Absorption (CPA). This is a phenomena characterized by all the energy injected into a system being completely absorbed, with no reflection or transmission [Chong2010, Fyodorov2017, Lei2020, Imani2020, Frazier2020, Hougne2021, Erb2024, Faul2025]. CPA has been found in optics [Gupta2012, Pichler2019, Baranov2017], acoustics [Song2014, Meng2017, IglesiasMartinez2025], heat transfer [Li2022], and quantum single photon systems [Vetlugin2021, Vetlugin2022_Resolution, Vetlugin2022_AHOM]. In microwave scattering, the signature of CPA is an eigenvalue of the $S$-matrix achieving complex zero, so $\text{det}S$ is the complex scalar function who’s zero is the enabling condition for CPA [Guo2023]. Alongside enabling CPA, $\text{det}S=0+i0$ also turns a two channel system into a tunable robust splitter [erb2025robustwavesplittersbased].
@@
Figure 1: Speckle pattern of scattering singularities in $d=2$ parameter space. The experimental S-matrix data comes from a tetrahedral microwave graph. (a-b) Magnitude of $\text{det}S$, red lines mark locations of $\text{Re}[\text{det}S]=0$ while black lines mark locations of $\text{Im}[\text{det}S]=0$. White symbols highlight intersections which correspond to $\text{det}S=0+i0$ which are CPAs. (c-d) Phase of $\text{det}S$, black circles show locations of phase windings which align exactly with the white symbols in (a-b). Arrows on the circles in (d) show the two nearby CPAs have opposite winding numbers.
Another singularity of interest is an Exceptional Point degeneracy of order $P$ (EP-P) where $P\geq 2$ eigenvalues and eigenvectors of a non-Hermitian operator become degenerate [Berry2004, El-Ganainy18, Alu19, Özdemir19, Ashida20, Ding2022]. Similar to CPAs, EPs have been explored in many fields such as optics [PhysRevLett.106.213901, PhysRevLett.112.203901, Hodaei2017, Chen2017], microwave cavities [PhysRevLett.86.787, PhysRevE.69.056216, PhysRevLett.106.150403], exciton-polaritons [Gao2015], acoustics [Shi2016], waveguides [Doppler2016], electrical circuits [Stehmann_2004, Suntharalingam2023], etc. In this paper we specifically consider EPs of the $S$-matrix, which in general are unrelated to the Hamiltonian EPs, except for specially contrived cases such as in Ref. [xia2025]. The EPs of the $S$-matrix are the same as those for the impedance $Z$, admittance $Y$, and Wigner reaction $K$ matrices [erb2025]. EP-2s are topologically protected as they are the zeros of the complex scalar field $\lambda_{j}-\lambda_{k}$ where $\lambda_{k}$ is the $k$th eigenvalue of an operator $\mathcal{H_{S}}$. Exceptional Points of order higher than $2$ are not topologically protected, as will be discussed later.
Due to the eigenvalue repulsion, the raw eigenvalue difference $\lambda_{j}-\lambda_{k}$ can actually be a poor quantity to use for identifying Exceptional Points. We instead use the eigenvector coalescence measure $|C_{jk}|$ where:
@@
[…]
Other simple singularities include matrix-element zeros of the $S$-matrix. The complex zeros of $S_{xx}$ or $S_{xy}$ are known as Reflectionless Scattering Modes (RSMs) and Transmissionless Scattering Modes (TSMs) [Sol2023, Davy2023, Jiang2024, Faul2025, Kang2021, Asano2016, Huang2022, Sweeney2020, Stone2021, Genack2024]. In a system with broken Lorentz reciprocity, it is possible to have $S_{xy}=0$ or $S_{yx}=0$ independently of each other, and when only one equals zero there is Uni-Directional Transmission. Further we can define $\delta R_{xy}:=S_{xx}-S_{yy}$ and $\delta T_{xy}:=S_{xy}-S_{yx}$ as the reflection and transmission differences between channels $x$ and $y$. If $\delta R_{xy}=0+i0$, then the two channels have symmetric reflections, which is an interesting singularity when it occurs in complex scattering systems without any geometric symmetries. Similarly if $\delta T_{xy}=0+i0$, then transmission between the two channels is reciprocal and we find that this singularity is present in systems with broken Lorentz reciprocity [erb2025].
As stated before, the divergence of complex time delays is associated with scattering singularities: Wigner-Smith $\tau_{WS}=\frac{-i}{M}\frac{\partial}{\partial E}\text{log}[\text{det}S]$ with CPA, Reflection $\tau_{xx}=-i\frac{\partial}{\partial E}\text{log}[S_{xx}]$ with RSM, Transmission $\tau_{xy}=-i\frac{\partial}{\partial E}\text{log}[S_{xy}]$ with TSM, etc. Since complex time delay itself can also take the value $0+i0$, there are additional scattering singularities that can be identified with complex time delay. To see why the zero of complex time delay is interesting, we turn to recent work on interpreting complex time delay in the frequency domain as the time and center frequency shift of a pulse in the time domain [Asano2016, giovannelli2025]. This implies that the zero of complex time delay corresponds to a minimal distortion point of the system, at which signals are only attenuated (or amplified), but not otherwise distorted after propagation.
There are a myriad of other scattering singularities that we could discuss, and a more expansive listing is provided in Appendix H.
@@ I.2 Higher Order Singularities
All the singularities given so far are topologically stable, being the zero of a single complex scalar field $\mathcal{S}$. In experimental systems, we can use embedded tunable perturbers to create collisions of fundamental singularities of different fields ($\mathcal{S}$ and $\mathcal{S}^{\prime}$) at real frequencies in a given parameter space. These collisions are not topologically protected and a small fluctuation in the system can move the singularities apart. Despite their instability, such “higher order” singularities can have novel properties and further applications, and so should be of great interest.
One example of a higher order singularity is the $S$-matrix CPA-EP demonstrated in Ref. [erb2025] to turn a complex, irregular tetrahedral graph into a robust 50:50 In-Phase/Quadrature microwave splitter. CPA-EPs have also been found in acoustic systems [xia2025], and in the spectrum of the Hamiltonian [Sweeney2019, Wang2021]. There are also higher order $S$-matrix EP-Ps which are intersections of exactly P1-1 unique EP-2s. Because the eigenvalue detuning around an EP has a $\frac{1}{P}$ power law, it shows extreme sensitivity to small perturbations, leading to proposals of EP sensors [PhysRevLett.112.203901, Hodaei2017, Chen2017, Kononchuk22]. RSMs and TSMs can also be made degenerate, for example if both $S_{xx}$ and $S_{yy}$ equal zero simultaneously then the system has an RSM-2, sometimes called a RSM or RL EP [Ferise2022] because the eigenvectors of a reflectionless operator become degenerate. RSM-2s have been used to create broadband perfect transmission [zhumabay2025]. Reciprocal systems have fundamental TSM-2s and can only have TSM-Ps of even order P.
@@
Again, there are many ways to combine fundamental singularities into higher order singularities, but none of these combinations are topologically stable and the more fundamental singularities included, the less probable such events become in a statistical ensemble.